A dumas Nitrogen analysis of 95.72mg sample of phenylene-diamine gave 1.3ml of Nitrogen at 20 degrees Celsius and 746mmHg.The gas was collected over saturated aqueous pottasium hydroxide solution ( let H2O be equal to 6mmHg) Calculate the percentage of Nitrogen in the compound
1
Expert's answer
2019-09-23T03:44:37-0400

n=(1.3/10^3*746/760*101.325)/(8.314*293.15)=0.00005304 mol

m= 0.00005304*28=0.001485 g

w=0.001485/0.09572*100%=1,55% ---- in sample!

but if you mean in compound, then it is 28/108*100%=26.9%

Answer: 1,55 % or 26.9%.